# Direct and Converse Theorems in Estimating Signals with Sublinear Sparsity

### v1

@@ III-B Hypothesis Testing
A hypothesis testing is considered. For some pdf $q(\boldsymbol{y})$ on $\mathbb{R}^{N}$, consider the null hypothesis $\boldsymbol{y}\sim q(\boldsymbol{y})$ for the received vector $\boldsymbol{y}$ in (1). The alternative hypothesis is the true marginal pdf $\boldsymbol{y}\sim p(\boldsymbol{y})$ of $\boldsymbol{y}$. Intuitively, it is impossible to estimate the signal vector $\boldsymbol{x}$ from the received vector $\boldsymbol{y}$ if the true pdf $p(\boldsymbol{y})$ is indistinguishable from $q(\boldsymbol{y})$. The difference between $p(\boldsymbol{y})$ and $q(\boldsymbol{y})$ is measured with the KL divergence.
@@
The condition $\sigma^{2}>u_{\mathrm{max}}^{2}/2$ cannot be weakened. Extreme value theory [39, p. 302] implies that the expected maximum of $N$ independent zero-mean Gaussian noise samples with variance $\sigma_{N/k}^{2}$ is $\sqrt{2\sigma_{N/k}^{2}\log N}\{1+o(1)\}=\sqrt{2\sigma^{2}(1-\gamma)^{-1}}$ $\cdot\{1+o(1)\}$. For $\sigma^{2}<u_{\mathrm{max}}^{2}/2$, thus, there is small $\gamma>0$ such that the maximum non-zero signal is strictly larger than all Gaussian noise samples. In this case, it is possible to detect the positions of the maximum non-zero signals from the observation $\boldsymbol{y}$.
@@ III-C Square Error
We prove a general lemma to connect the KL divergence $D(p\|q)$ with the square error. Consider a general channel that is determined with any conditional pdf $p_{\boldsymbol{y}|\boldsymbol{x}}$ of $\boldsymbol{y}\in\mathbb{R}^{N}$ given $\boldsymbol{x}\in\mathbb{R}^{N}$. We write the capacity of this channel—the supremum of the mutual information $I(\boldsymbol{x};\boldsymbol{y})$ over the distribution of $\boldsymbol{x}$ under the average power constraint $k^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]\leq|\mathcal{M}|^{-1}\sum_{m\in\mathcal{M}}u_{m}^{2}$—as $C(p_{\boldsymbol{y}|\boldsymbol{x}})$. For any pdf $q(\boldsymbol{y})$ of $\boldsymbol{y}$, define
$J(q(\boldsymbol{y}))=C(p_{\boldsymbol{y}|\boldsymbol{x}})-\mathbb{E}\left[\log\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})}{q(\boldsymbol{y})}\right]+D(p(\boldsymbol{y})\|q(\boldsymbol{y})).$ (28)
It is straightforward to confirm $J(q(\boldsymbol{y}))\geq 0$ from the definition of the capacity $C(p_{\boldsymbol{y}|\boldsymbol{x}})$ and the positivity of the KL divergence in (24). The following lemma is proved for a general channel $p_{\boldsymbol{y}|\boldsymbol{x}}$ that contains the AWGN channel (1).
@@
Let $\hat{\boldsymbol{x}}(\boldsymbol{y})\in\mathbb{R}^{N}$ denote any estimator of $\boldsymbol{x}$ given $\boldsymbol{y}$. Then, for any pdf $q(\boldsymbol{y})$ we have
$\frac{1}{k}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq\frac{1}{\alpha}\left\{C_{N}(\alpha)e^{-2J(q(\boldsymbol{y}))}-1\right\}$ (29)
@@
$C_{N}(\alpha)=\frac{(1-u_{\mathrm{max}}^{4}\alpha^{2})e^{\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}}{(1+f(\alpha)\{[1+(N/k-1)^{-1}]^{k}-1\})^{2}},$ (30)
@@
See Appendix E. ∎
Lemma 3 only provides a trivial bound when $J(q(\boldsymbol{y}))$ tends to infinity in the sublinear sparsity limit for all $q(\boldsymbol{y})$. In this case, the lower bound (29) reduces to $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq-\alpha^{-1}$ in the sublinear sparsity limit. When there is some pdf $q(\boldsymbol{y})$ such that $J(q(\boldsymbol{y}))\to 0$ holds in the sublinear sparsity limit, on the other hand, we have
@@
The condition $J(q(\boldsymbol{y}))\to 0$ in (28) implies that $q(\boldsymbol{y})$ is the capacity-achieving output pdf and indistinguishable from the true output pdf $p(\boldsymbol{y})$ in terms of the KL divergence: $C(p_{\boldsymbol{y}|\boldsymbol{x}})-\mathbb{E}[\log p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})/q(\boldsymbol{y})]\to 0$ and $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))\to 0$.
@@
Let $q(\boldsymbol{y})$ denote the zero-mean Gaussian pdf with covariance $N^{-1}\mathbb{E}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$. Lemma 2 implies $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))\to 0$ in the sublinear sparsity limit under the assumptions in Theorem 2. Thus, it is sufficient to confirm that this pdf $q(\boldsymbol{y})$ is the capacity-achieving output pdf for the AWGN channel (1).
We evaluate the second term in (28). From the definitions of $q(\boldsymbol{y})$ and the AWGN channel in (1) we have
$\mathbb{E}\left[\log\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})}{q(\boldsymbol{y})}\right]$
$=\frac{N}{2}\log\frac{N^{-1}\mathbb{E}[\|\boldsymbol{y}\|_{2}^{2}]}{\sigma_{N/k}^{2}}-\frac{\mathbb{E}[\|\boldsymbol{y}-\boldsymbol{x}\|_{2}^{2}]}{2\sigma_{N/k}^{2}}+\frac{N}{2}$
$=\frac{N}{2}\log\left(1+\frac{\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}{N\sigma_{N/k}^{2}}\right)=C(p_{\boldsymbol{y}|\boldsymbol{x}}),$ (33)
where the second equality follows from the definition of the AWGN channel in (1) while the last equality is the definition itself for the capacity of the AWGN channel under the power constraint $k^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]\leq|\mathcal{M}|^{-1}\sum_{m\in\mathcal{M}}u_{m}^{2}$. Thus, Theorem 2 holds. ∎
@@ IV Numerical Results
IV-A Non-Separable Bayesian Estimator
As shown shortly, the ML estimator (6) in Proposition 1 outperforms the separable Bayesian estimator [1] in the low noise regime. However, it is inferior to the separable Bayesian estimator in the high noise regime. To improve the performance of the ML estimator in the high noise regime, this paper proposes the following non-separable Bayesian estimator of $x_{n}$:
$\hat{x}_{n}=\frac{|\mathcal{M}|^{-1}\sum_{m\in\mathcal{M}}u_{m}e^{u_{m}y_{n}/\sigma_{N/k}^{2}-\ell_{m}(\boldsymbol{y}_{\setminus n})/(2\sigma_{N/k}^{2})}}{1+|\mathcal{M}|^{-1}\sum_{m\in\mathcal{M}}e^{u_{m}y_{n}/\sigma_{N/k}^{2}-\ell_{m}(\boldsymbol{y}_{\setminus n})/(2\sigma_{N/k}^{2})}},$ (34)
@@ IV-A Non-Separable Bayesian Estimator
with $\xi_{k}^{N}(k_{0};\boldsymbol{y})$ given in (9). See Appendix A-B for the derivation of (34).
Figure 1: Square error versus SNR $1/\sigma^{2}$ in dB for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, and the AWGN channel (1). $10^{5}$ independent trials were simulated for the non-separable Bayesian estimator (34), separable Bayesian estimator [1], and ML estimator (6). The vertical dashed line shows the threshold $1/\sigma^{2}=2$ in Theorems 1 and 2.
The constant non-zero signals $\mathcal{U}=\{1\}$ were considered in numerical simulations to compare the three estimators for the AWGN channel (1). Figure 1 shows the square error normalized by $k$ versus SNR $1/\sigma^{2}$. The ML estimator outperforms the separable Bayesian estimator [1] in the high SNR regime while it is inferior to the separable Bayesian estimator in the low SNR regime. The non-separable Bayesian estimator (34) outperforms the ML estimator (6) for all SNRs. However, it is inferior to the separable Bayesian estimator in the low SNR regime. The high and low SNR regimes are separated at approximately 6 dB. The suboptimality of the non-separable Bayesian estimator is due to the max-log approximation in the high SNR regime.
@@ IV-B Compressed Sensing
The separable and non-separable Bayesian estimators are applied to AMP [1] for compressed sensing with sublinear sparsity. The ML estimator is not considered, because of its non-differentiability. In compressed sensing, the $M$-dimensional measurement vector $\boldsymbol{y}\in\mathbb{R}^{M}$ is given by
@@
Figure 2: Square errors of AMP using separable and non-separable Bayesian estimators versus $\delta=M/\{k\log(N/k)\}$ for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, $1/\sigma^{2}=40$ dB, 3030 iterations, and the measurement model (36). According to the numerical simulations shown in Fig. 1, in non-separable AMP, the non-separable Bayesian estimator was used when the input SNR to the denoiser is larger than 6 dB. Otherwise, the separable Bayesian estimator was used. $10^{5}$ independent trials were simulated. The optimal threshold [12] is approximately $\delta_{\mathrm{opt}}\approx 0.5$ while the threshold for AMP [1] is $\delta_{\mathrm{AMP}}\approx 2$.
Figure 2 shows the square errors for AMP using the separable and non-separable Bayesian estimators. The square errors reduce rapidly as $\delta$ increases toward the threshold $\delta_{\mathrm{AMP}}\approx 2$ for AMP [1]. In contrast to the superiority of the non-separable Bayesian estimator in the high SNR regime, the two AMP algorithms are comparable to each other for all $\delta$.
To understand this unexpected result, we focus on the input-output relationship for denoising shown in Fig. 3. The iteration trajectories for AMP run from the upper right toward the lower left. If perfect Gaussianity held for estimation errors before denoising in AMP, the two trajectories for AMP would overlap the corresponding input-output relationship for the separable/non-separable Bayesian estimators. However, AMP outputs much worse square errors than the corresponding Bayesian estimator for small input square errors while it overlaps the separable Bayesian estimator for large input square errors. Furthermore, AMP using the non-separable estimator is trapped in worse input square errors than the separable counterpart on the final stage of convergence. As a result, the two AMP algorithms achieve almost the same performance.
These observations imply that the performance bottleneck is in non-ideal properties for estimation errors before denoising, rather than the type of denoisers. Because of finite size effect, the estimation errors cannot be regarded as zero-mean i.i.d. Gaussian random variables. This non-Gaussianity is a dominant factor for the degradation of the convergence property in finite-sized systems.
Figure 3: Square errors after denoising versus the square error normalized by NN before denoising for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, and $1/\sigma^{2}=40$ dB. The measurement model (36) was postulated for the separable/non-separable AMP while the AWGN channel (1) was assumed for the separable/non-separable Bayesian estimators. In non-separable AMP, the non-separable Bayesian estimator was used when the input SNR to the denoiser is larger than 6 dB. Otherwise, the separable Bayesian estimator was used. $10^{5}$ independent trials were simulated.

### v3

@@ III-B Hypothesis Testing
A hypothesis testing is considered. For some pdf $q(\boldsymbol{y})$ on $\mathbb{R}^{N}$ different from the true marginal pdf $p(\boldsymbol{y})$ of $\boldsymbol{y}$, consider the null hypothesis $\boldsymbol{y}\sim q(\boldsymbol{y})$ for the received vector $\boldsymbol{y}$ in (1). The alternative hypothesis is the true marginal pdf $\boldsymbol{y}\sim p(\boldsymbol{y})$. Intuitively, it is impossible to estimate the signal vector $\boldsymbol{x}$ from the received vector $\boldsymbol{y}$ if the true pdf $p(\boldsymbol{y})$ is asymptotically indistinguishable from $q(\boldsymbol{y})$. The difference between $p(\boldsymbol{y})$ and $q(\boldsymbol{y})$ is measured with the KL divergence.
@@
The condition $\sigma^{2}>u_{\mathrm{max}}^{2}/2$ cannot be weakened. Extreme value theory [40, p. 302] implies that the expected maximum of $N$ independent zero-mean Gaussian noise samples with variance $\sigma_{N/k}^{2}$ is $\sqrt{2\sigma_{N/k}^{2}\log N}\{1+o(1)\}=\sqrt{2\sigma^{2}(1-\gamma)^{-1}}$ $\cdot\{1+o(1)\}$. For $\sigma^{2}<u_{\mathrm{max}}^{2}/2$, thus, there is small $\gamma>0$ such that the maximum non-zero signal is strictly larger than all Gaussian noise samples. In this case, it is possible to detect the positions of the maximum non-zero signals from the observation $\boldsymbol{y}$.
@@ III-C Square Error
We prove a lemma to connect the KL divergence $D(p\|q)$ with the square error. For notational convenience, we write the conditional pdf $p(\boldsymbol{y}|\boldsymbol{x})$ representing the AWGN channel (1) as $p_{\boldsymbol{y}|\boldsymbol{x}}$. This notation is useful in utilizing the permutation invariance of pdfs. We know that the capacity of the AWGN channel (1)—the supremum of the mutual information $I(\boldsymbol{x};\boldsymbol{y})$ over the distribution of $\boldsymbol{x}$ under the average power constraint $N^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]$—is given by
$C_{\mathrm{AWGN}}=\frac{1}{2}\log\left(1+\frac{\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}{N\sigma_{N/k}^{2}}\right),$ (28)
which is achieved by $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{0},N^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]\boldsymbol{I}_{N})$. In this case, the capacity-achieving output pdf is equal to the zero-mean Gaussian pdf with covariance $N^{-1}\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$, i.e. $q(\boldsymbol{y})$ in Lemma 2.
For this Gaussian pdf $q(\boldsymbol{y})$, define
$J(q(\boldsymbol{y}))$ $=C_{\mathrm{AWGN}}-\mathbb{E}\left[\log\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})}{q(\boldsymbol{y})}\right]+D(p(\boldsymbol{y})\|q(\boldsymbol{y}))$
$=D(p(\boldsymbol{y})\|q(\boldsymbol{y})),$ (29)
where the second equality holds because $q(\boldsymbol{y})$ is equal to the capacity-achieving output pdf for the AWGN channel (1). The following lemma is proved for the capacity-achieving output pdf $q(\boldsymbol{y})$ and the AWGN channel (1), while it is possible to generalize $q(\boldsymbol{y})$ and the conditional pdf $p_{\boldsymbol{y}|\boldsymbol{x}}$ to any pdf and any permutation-invariant conditional pdf, i.e. $p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{P}\boldsymbol{y}|\boldsymbol{P}\boldsymbol{x})=p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})$ for any permutation matrix $\boldsymbol{P}$.
@@
Let $\hat{\boldsymbol{x}}(\boldsymbol{y})\in\mathbb{R}^{N}$ denote any estimator of $\boldsymbol{x}$ given $\boldsymbol{y}$ in the AWGN channel (1). Then, for the zero-mean Gaussian pdf $q(\boldsymbol{y})$ with covariance $N^{-1}\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$ we have
$\frac{1}{k}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq\frac{1}{\alpha}\left\{C_{N}(\alpha)e^{-2D(p(\boldsymbol{y})\|q(\boldsymbol{y}))}-1\right\}$ (30)
@@
$C_{N}(\alpha)=\frac{(1-u_{\mathrm{max}}^{4}\alpha^{2})e^{\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}}{\{1+f(\alpha)[(1-k/N)^{-k}-1]\}^{2}},$ (31)
@@
See Appendix G. ∎
Lemma 3 only provides a trivial bound when $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))$ tends to infinity in the sublinear sparsity limit. In this case, the lower bound (30) reduces to $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq-\alpha^{-1}$ in the sublinear sparsity limit. When $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))\to 0$ holds in the sublinear sparsity limit, on the other hand, we have
@@
Let $q(\boldsymbol{y})$ denote the zero-mean Gaussian pdf with covariance $N^{-1}\mathbb{E}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$. Lemma 2 implies $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))\to 0$ in the sublinear sparsity limit under the assumptions in Theorem 2. Under the additional assumption $\gamma<1/2$, Lemma 3 implies Theorem 2. Thus, it is sufficient to confirm that the assumption (3) in Theorem 2 implies $\gamma<1/2$.
Under the change of variables $u=u_{\mathrm{max}}/\sigma$ and $v=u_{\mathrm{min}}/\sigma$, we write the first argument of the minimum operator in the assumption (3) as
$\gamma_{1}(u,v)=\frac{C_{0}}{C_{0}+(uv)^{-1}},\quad C_{0}=\frac{1}{u^{2}}\left(1-\frac{u^{2}}{2}\right).$ (34)
We prove $\gamma_{1}(u,v)<1/2$ for all $u\in(0,\sqrt{2})$ and $v\in(0,u]$. From this inequality, we find that the assumption (3) implies $\gamma<1/2$. Since $\gamma_{1}(u,v)$ is monotonically increasing with respect to $v\in(0,u]$ for fixed $u\in(0,\sqrt{2})$, we obtain $\gamma_{1}(u,v)\leq\gamma_{1}(u,u)\equiv\gamma_{1}(u)$, given by
$\gamma_{1}(u)=\frac{C_{0}u^{2}}{C_{0}u^{2}+1}=\frac{2-u^{2}}{4-u^{2}}<\frac{1}{2}$ (35)
for all $u^{2}\in(0,2)$. Thus, $\gamma_{1}(u,v)<1/2$ holds for all $u\in(0,\sqrt{2})$ and $v\in(0,u]$. ∎
@@ IV Numerical Results
IV-A Non-Separable Approximate Bayesian Estimator
As shown shortly, the ML estimator (6) in Proposition 1 outperforms the separable Bayesian estimator [1] in the low noise regime. However, it is inferior to the separable Bayesian estimator in the high noise regime. To improve the performance of the ML estimator in the high noise regime, we propose the following non-separable approximate Bayesian estimator of $x_{n}$—simply called non-separable estimator:
$\hat{x}_{n}=\frac{\sum_{m\in\mathcal{M}}u_{m}e^{u_{m}y_{n}/\sigma_{N/k}^{2}-\ell_{m}(\boldsymbol{y}_{\setminus n})/(2\sigma_{N/k}^{2})}}{1+\sum_{m\in\mathcal{M}}e^{u_{m}y_{n}/\sigma_{N/k}^{2}-\ell_{m}(\boldsymbol{y}_{\setminus n})/(2\sigma_{N/k}^{2})}},$ (36)
@@ IV-A Non-Separable Approximate Bayesian Estimator
with $\xi_{k}^{N}(k_{0};\boldsymbol{y})$ given in (9). See Appendix A-C for a heuristic derivation of (36).
Figure 2: Square error versus SNR $1/\sigma^{2}$ in dB for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, and the AWGN channel (1). $10^{5}$ independent trials were simulated for the non-separable estimator (36), separable Bayesian estimator [1], and ML estimator (6). The vertical dashed line shows the threshold $1/\sigma^{2}=2$ in Theorems 1 and 2.
The constant non-zero signals $\mathcal{U}=\{1\}$ were considered in numerical simulations to compare the three estimators for the AWGN channel (1). Figure 2 shows the square error normalized by $k$ versus SNR $1/\sigma^{2}$. The ML estimator outperforms the separable Bayesian estimator [1] in the high SNR regime while it is inferior to the separable Bayesian estimator in the low SNR regime. The non-separable estimator (36) outperforms the ML estimator (6) for all SNRs. However, it is inferior to the separable Bayesian estimator in the low SNR regime. The high and low SNR regimes are separated at approximately 6 dB. The suboptimality of the non-separable estimator (36) is due to the max-log approximation in the high SNR regime.
@@ IV-B Compressed Sensing
The separable Bayesian estimator and the non-separable estimator (36) are applied to AMP [1] for compressed sensing with sublinear sparsity. The ML estimator is not considered, because of its non-differentiability. In compressed sensing, the $M$-dimensional measurement vector $\boldsymbol{y}\in\mathbb{R}^{M}$ is given by
@@
AMP using the separable Bayesian estimator was proved to achieve vanishing square error in the sublinear sparsity limit [1, Proposition 1] if $\delta$ is larger than the threshold $\delta_{\mathrm{AMP}}=2(1+\sigma^{2})$. This threshold $\delta_{\mathrm{AMP}}$ is strictly larger than the optimal threshold [13] $\delta_{\mathrm{opt}}=2/\log(1+\sigma^{-2})$.
@@
Figure 3: Square errors of AMP using the separable Bayesian estimator and the non-separable estimator versus $\delta=M/\{k\log(N/k)\}$ for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, $1/\sigma^{2}=40$ dB, 3030 iterations, and the measurement model (38). According to the numerical simulations shown in Fig. 2, in non-separable AMP, the non-separable estimator (36) was used when the input SNR to the denoiser is larger than 6 dB. Otherwise, the separable Bayesian estimator was used. $10^{5}$ independent trials were simulated. The optimal threshold [13] is approximately $\delta_{\mathrm{opt}}\approx 0.22$ while the threshold for AMP [1] is $\delta_{\mathrm{AMP}}\approx 2.0$.
Figure 3 shows the square errors for AMP using the separable Bayesian estimator and the non-separable estimator (36). The square errors reduce rapidly as $\delta$ increases toward the threshold $\delta_{\mathrm{AMP}}\approx 2$ for AMP [1]. In contrast to the superiority of the non-separable estimator in the high SNR regime, the two AMP algorithms are comparable to each other for all $\delta$.
To understand this unexpected result, we focus on the input-output relationship for denoising shown in Fig. 4. The iteration trajectories for AMP run from the upper right toward the lower left. If perfect Gaussianity held for estimation errors before denoising in AMP, the two trajectories for AMP would overlap the corresponding input-output relationship for the separable Bayesian estimator and the non-separable estimator. However, AMP outputs much worse square errors than the corresponding estimators for small input square errors while it overlaps the separable Bayesian estimator for large input square errors. Furthermore, AMP using the non-separable estimator is trapped in worse input square errors than the separable counterpart on the final stage of convergence. As a result, the two AMP algorithms achieve almost the same performance.
These observations suggest that the performance bottleneck is in non-ideal properties for estimation errors before denoising, rather than the type of denoisers. Because of finite size effect, the estimation errors cannot be regarded as zero-mean i.i.d. Gaussian random variables. This non-Gaussianity is a dominant factor for the degradation of the convergence property in finite-sized systems.
Figure 4: Square errors after denoising versus the square error normalized by NN before denoising for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, and $1/\sigma^{2}=40$ dB. The measurement model (38) was postulated for the separable/non-separable AMP while the AWGN channel (1) was assumed for the separable Bayesian/non-separable estimators. In non-separable AMP, the non-separable estimator (36) was used when the input SNR to the denoiser is larger than 6 dB. Otherwise, the separable Bayesian estimator was used. $10^{5}$ independent trials were simulated.